Need help asap

Qvestion & (of 7 6. points Problem 2-15 Using Income Statements [LO 2] o retained eamings $3075 dvidends paid $1,020 intenestexpense $1.300:tax ra30 percent Caloulate the depreciaion expense for the company. (Do not round intermediate calculations and Problem 2-15 Using ncome from cash ow LO 21 MacBook Pro 0 4 6
Answer #1

Please provide rating..

Computation of EBITDA
a Addition to retained earning = 3075
b Dividend paid = 1020
c=a+b Net income 4095
d=c/70% Earning before tax = 5850
e Interest expenses 1390
f=d+e EBITDA 7240
Computation of gross margin
a Sales 52900
b Cost 37900
c=a-b Gross margin 15000
Computation of depreciation
Depreciation = Gross margin - EBITDA
=15000-7240 7760
Depreciation = 7760